OpenStudy (anonymous):

A ball is dropped from the top of a 286foot building.How many seconds will it take to hit the ground if the height is given by h = −16t2 + s,where (s) is the initial height?

OpenStudy (anonymous):

when the ball hits the ground the height is zero: 0 = -16t^2 + 286 16t^2 = 286 t^2 = 286/16 \[t = \sqrt{286/16}\] t =about 4.2s

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

no problem